OpenStudy (anonymous):

A mega-super-really-quick question? In Algebra, when I solve, I always get something that = x. Ex: 2x=0 x=-2 BUT, now in a specific problem, I don't have an x, and I'm still expected to solve. Ex: 0=6 In this scenario, what do I do?

OpenStudy (anonymous):

first of all , 2x=0 is not x=-2 second of all, there is no way that 0=6. There is nothing you can say in this case

OpenStudy (anonymous):

I'm talking about when graphing rational functions. One of my problems: 2/x^2-4 My teachers says set it = to zero and cancel out the bottom. Sorry about the example up top. I was trying to think of a fast one. :/

OpenStudy (anonymous):

I'm trying to find the x-intercept etc.

OpenStudy (anonymous):

2/x^2=4 x^2=1/2 x=sqrt (1/2)

OpenStudy (mertsj):

Factor the denominator and set it to 0. You will find that x = 2 or -2. Those numbers cause the denominator to be 0 so x cannot be those numbers. These are vertical asymptotes: x=2 and x = -2

OpenStudy (mertsj):

There are no x intercepts because the numerator can never be 0 so the fraction can never be 0. If the fraction cannot be 0, y cannot be 0 and if y is not 0 there is no x intercept.

OpenStudy (anonymous):

Okay, thanks. I was trying to cancel out the denominator instead.